OpenStudy (anonymous):

Fan and medal A rectangular meatloaf is 8 inches long, 5 inches wide, and 4 inches tall. What is the volume of the meatloaf? Select the best answer from the choices provided. 320 in3 256 in3 160 in3 100 in3 Give explanation .

OpenStudy (mathmath333):

V=l*b*h=8*5*4

OpenStudy (anonymous):

?